Last year, 46% of business owners gave a holiday gift to their employees. A survey of business owners indicated that 50% plan to provide a holiday gift to their employees. Suppose the survey results are based on a sample of 120 business owners

use this sample information to test whether the proportion of business owners providing a holiday gift has increased. use the .10 level of significance. round critical value 2 decimals, round test statistic 2 decimal
Answer #1

H0: p = 0.46 ,

Ha: p > 0.46

Test statistics

z = (\hat{p} - p) / sqrt [ p ( 1 - [) / n]

= ( 0.50 - 0.46 ) / sqrt ( 0.46 ( 1 - 0.46) / 120)

= 0.88

This is test statistics value.

p-value = P(Z > z) = P(Z > 0.88 )

= 0.1894

Since p-value > 0.10 level, do not reject H0.

We conclude at 0.10 level that we fail to support the claim